OpenStudy (anonymous):

find the maximum y value on the graph of y= f(x). f(x)= -x^2 +6x+10

OpenStudy (anonymous):

well that was wrong too many minus signs

OpenStudy (anonymous):

max is when \(x=-\frac{b}{2a}=-\frac{6}{2\times (-1)}=3\) replace \(x\) by \(3\) get \[y=-3^2+6\times 3+10=-9+18+10=9+10=19\]